If log20 = x then, Find the exact value of B such that log B = x + 3
Answer by jsmallt9(3758)

log(20) = x
log(B) = x + 3
One solution:
Substitute for x:
log(B) = log(20) + 3
Subtract log(20):
log(B) - log(20) = 3
Using a property of logarithms, log(p) - log(q) = log(p/q), to combine the two logs:
log(B/20) = 3
Rewrite in exponential form. In general, log(p) = n is equivalent to p = 10^n. Using this pattern on our equation we get:
B/20 = 10^3
which simplifies to:
B/20 = 1000
Multiply by 20:
B = 20000